Part D and E please
As in Problem 1, once again, you are borrowing $6500 to purchase a car. However, now the first payment is due immediately. There will be a total of 36 monthly payments (The first payment occurs immediately. The remaining 35 occur at the end of each of the following months). The advertised interest rate is 9% APR (therefore use a monthly discount rate of r-0.75% in your computations). The payments will all be equal in size. HINT: This is an annuity due problem. Make sure you understand the discussion in the book, especially Example 5B- paying attention to Footnote 3 (which applies to Parts B, C, and D). Also work the Annuity Due problems at the end of the chapter before you work this problem.
Part D) Now assume that the size of the car payments decrease by 0.75% every month. What is the size of the payment that occurs 15 months from today? Part E)Now assume that the size of the car payments increases by 0.75% every month. What is the size of the payment that occurs 15 months from today? Now repeat Problem 2-B (in which you borrow S6500, make a total of 36 payments with the first
Answer #1

Answer:

Part D)

Let x be the initial payment

So g=-0.75% every month

r=0.75%

n=35 months

P= $6500

P=x*(1+r)*(1-{(1+g)/(1+r)}^36)/(r-g)

6500=X*(1+0.75%)*(1-{(1-0.75%)/(1+0.75%)}^36)/(0.75%-(-0.75%))

X=$231.92

So payment occurs from 15 month today= 231.92*(1-0.75%)^15=$207.163

Part E)

Let x be the initial payment

So g=0.75% every month

r=0.75%

n=35 months

P= $6500

P=x+x*(1+g)*n/(1+i)

6500=x+x*(1+0.75%)*35/(1+0.75%)

X=180.55

So payment occurs from 15 month today= 180.55*(1+0.75%)^15=$201.97